OpenStudy (anonymous):

Name the property of equality that justifies: If AB= CD and CD= EF than AB= EF A. Reflexive B. Symmetric C. Transitive D. Substitution

OpenStudy (anonymous):

c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Name the property of equality that justifies: If x+3=5 than x=2 . A. Multiplication Property of Equality B. Division Property of Equality C. Distributive Property of Equality D. Subtraction Property of Equality

OpenStudy (anonymous):

c transitive property.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

u wrong bro

OpenStudy (anonymous):

Name the property of equality that justifies: If x=3 and y=x+5 then y=8 A. Addition Property of Equality B. Distributive Propery of Equality C. Substitution Property of Equality D. Reflexive Property of Equality

OpenStudy (anonymous):

this is functional dependency from dbms.

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

C would be correct :)